# A Survey on Diffusion Models for Anomaly Detection

### v2

@@ 5 Discussions
Computational Cost
Widespread adoption of DMAD remains constrained by substantial computational requirements, particularly when processing high-dimensional data or extended time series. Computational demands manifest during both distribution learning in training and data generation in sampling phases, with the sampling process’s iterative nature demanding numerous steps for quality output Cui et al. (2023). For example, high-resolution IAD suffers from increased computational overhead due to data volume, model complexity, and the need to capture temporal dependencies Wyatt et al. (2022).  However, some researchers are actively exploring solutions.  One promising direction is faster sampling methods, such as progressive distillation and optimized sampling schedules like align your steps, which aim to reduce sampling steps while maintaining quality Salimans and Ho (2021).  Another approach is model compression through techniques like pruning and quantization.  Additionally, efficient architectures, like Cui et al. (2023) leveraging ER SDEs for faster sampling, and preconditioning methods offer further computational gains.
Interpretability and Explainability
Interpretability remains a key challenge for DMAD Katsuoka et al. (2024).  Understanding DM’s decision-making is crucial, especially in critical applications.  Visualizing anomaly scores, like spatially highlighting anomalous image regions as in reconstruction-based methods Yao et al. (2025), becomes essential.  However, the iterative denoising process inherent in DMs complicates interpretability.  Explaining anomaly detection requires identifying and explaining deviating features.  Integrating DMs with explainable AI (xAI) techniques, such as incorporating attention mechanisms or leveraging LLMs for textual/visual explanations Wang et al. (2023b), offers a promising research direction.
Complex Data Distributions
@@
Robustness and Adversarial Attacks
Adversarial robustness is a critical concern for DMAD.  Similar to other deep learning models, DMs are vulnerable to adversarial perturbations, raising concerns about the reliability of diffusion-based AD systems Chen et al. (2024b).  For example, Kang et al. (2025) demonstrates the potential of DMs for adversarial defense in classification, directly applying such methods to AD can reduce anomaly detection rates. Specifically, the purification process may remove crucial anomaly signals along with noise.  Additionally, adversarial examples exhibit misalignment within DM manifolds, offering a potential detection avenue but requiring further investigation.  The observed robustness differences between pixel-space diffusion models (PDMs) and more vulnerable latent diffusion models (LDMs) Graham et al. (2023) further underscore the need to consider specific DM types.  Consequently, future research should prioritize robust DM-based AD methods, including architectures and training procedures that distinguish genuine anomalies from adversarial noise.  Promising directions include defense strategies like DIFFender Kang et al. (2025), leveraging text-guided diffusion and the adversarial anomaly perception phenomenon, and incorporating insights from adversarial example behavior within DM manifolds.
Edge-Cloud Collaboration
Real-time DMAD faces significant adoption barriers due to their substantial computational demands Li et al. (2023). Edge-cloud collaboration offers a promising solution to this challenge by distributing the workload between edge devices and cloud servers. Lightweight DMs deployed on edge devices perform initial anomaly screening, while resource-intensive tasks like full reconstructions are efficiently offloaded to cloud servers Yan et al. (2024). In addition, federated learning enables collaborative model training across edge devices and the cloud without sharing sensitive data, thereby enhancing generalization and preserving privacy Liu et al. (2024b). Dynamic DM partitioning Chen et al. (2024a) facilitates adaptive resource allocation, optimizing performance based on network conditions and computational demands. Strategic data placement and distributed DNN deployment principles further enhance system efficiency. The integration of proactive detection mechanisms, exemplified by Maat Lee et al. (2023b), proves particularly valuable for time-critical applications in cloud monitoring and AIOps environments.
Integrating with Large Language Models
@@ 6 Conclusion
In this survey, we introduce the core concepts of anomaly detection and diffusion models, providing a foundation for understanding DMAD. We systematically reviewed existing methodologies and their tasks across diverse data types, including image, time series, and multimodal data. Furthermore, we analyzed representative approaches for each data type, highlighting their strengths and limitations. To promote further research, we identified key challenges and promising future directions, aiming to advance DMAD research and inspire innovative solutions for real-world applications.

### v5

@@ 5 Discussions
Computational Cost.
Widespread adoption of DMAD remains constrained by substantial computational requirements, particularly when processing high-dimensional data or extended time series. Computational demands manifest during both distribution learning in training and data generation in sampling phases, with the sampling process’s iterative nature demanding numerous steps for quality output Cui et al. (2023). For example, high-resolution IAD suffers from increased computational overhead due to data volume, model complexity, and the need to capture temporal dependencies Wyatt et al. (2022); Sui et al. (2024).  However, some researchers are actively exploring solutions.  One promising direction is faster sampling methods, such as progressive distillation and optimized sampling schedules like align your steps, which aim to reduce sampling steps while maintaining quality Salimans and Ho (2021).  Another approach is model compression through techniques like pruning and quantization.  Additionally, efficient architectures, like Cui et al. (2023) leveraging ER SDEs for faster sampling, and preconditioning methods offer further computational gains.
Interpretability and Explainability.
Interpretability remains a key challenge for DMAD Katsuoka et al. (2024).  Understanding DM’s decision-making is crucial, especially in critical applications Naval Marimont et al. (2024).  Visualizing anomaly scores, like spatially highlighting anomalous image regions as in reconstruction-based methods Yao et al. (2025), becomes essential.  However, the iterative denoising process inherent in DMs complicates interpretability.  Explaining anomaly detection requires identifying and explaining deviating features.  Integrating DMs with explainable AI (xAI) techniques, such as incorporating attention mechanisms or leveraging LLMs for textual/visual explanations Wang et al. (2023b), offers a promising research direction.
Complex Data Distributions.
@@
Robustness and Adversarial Attacks.
Adversarial robustness is a critical concern for DMAD.  Similar to other deep learning models, DMs are vulnerable to adversarial perturbations, raising concerns about the reliability of diffusion-based AD systems Chen et al. (2024b).  For example, Kang et al. (2025) demonstrates the potential of DMs for adversarial defense in classification, directly applying such methods to AD can reduce anomaly detection rates. Specifically, the purification process may remove crucial anomaly signals along with noise.  Additionally, adversarial examples exhibit misalignment within DM manifolds, offering a potential detection avenue but requiring further investigation.  The observed robustness differences between pixel-space diffusion models (PDMs) and more vulnerable latent diffusion models (LDMs) Graham et al. (2023) further underscore the need to consider specific DM types.  Consequently, future research should prioritize robust DM-based AD methods, including architectures and training procedures that distinguish genuine anomalies from adversarial noise.  Some promising directions include defense strategies like DIFFender Kang et al. (2025), leveraging text-guided diffusion and the adversarial anomaly perception phenomenon, and incorporating insights from adversarial example behavior within DM manifolds.
Edge-Cloud Collaboration.
Real-time DMAD faces significant adoption barriers due to their substantial computational demands Li et al. (2023). Edge-cloud collaboration offers a promising solution to this challenge by distributing the workload between edge devices and cloud servers. Carefully designed Lightweight DMs deployed on edge devices perform initial anomaly screening, while resource-intensive tasks like full reconstructions are efficiently offloaded to cloud servers Yan et al. (2024). In addition, federated learning enables collaborative model training across edge devices and the cloud without sharing sensitive data, thereby enhancing generalization and preserving privacy Jin et al. (2024). Dynamic DM partitioning Chen et al. (2024a) facilitates adaptive resource allocation, optimizing performance based on network conditions and computational demands. Strategic data placement and distributed DNN deployment principles further enhance system efficiency. The integration of proactive detection mechanisms, exemplified by Maat Lee et al. (2023b), proves particularly valuable for time-critical applications in cloud monitoring and AIOps environments.
Integrating with Large Language Models.
@@ 6 Conclusion
In this survey, we first introduce the core concepts of anomaly detection and diffusion models, providing a foundation for understanding DMAD. Then, we systematically review existing methodologies and their tasks across diverse data types, including image, time series, and multimodal data. Furthermore, we analyze representative approaches for each data type, highlighting their strengths and limitations. To promote further research, we identify key challenges and promising future directions, aiming to advance DMAD research and inspire innovative solutions for real-world applications.